OpenStudy (anonymous):

O is any point in the plane of the square ABCD. The vectors OA and OB and OC are a,b and c respectively. Find the vector OD interms of a,b,c.

OpenStudy (anonymous):

a+c-b?

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

find BA OD=OC+CD=OC+BA

OpenStudy (anonymous):

OB=OA+AB AB=OB-OA BA=OA-OB

OpenStudy (anonymous):

THANK YOU VERY MUCH